OpenStudy (anonymous):

What are two number that add to 2 but multiply to negative one? One or both of them may be imaginary

OpenStudy (anonymous):

a + b = 2 a = 2 - b ab = -1 b(2-b) = -1 -b^2+ 2b = -1 0 = b^2 - 2b -1 solve this quadratic equation to find b

OpenStudy (anonymous):

Im solving it

OpenStudy (anonymous):

b = 2 +/- sqrt(4-4(-1) / 2 b = \[\frac{2\pm \sqrt{8}}{2} = 1 \pm \sqrt{2}\]

OpenStudy (anonymous):

Thank you, i understand it now, I was thinking of the wrong thing.